# One missing piece in Vision and Language: A Survey on Comics Understanding

### v1

@@ 7.1.2 Character Re-Identification
Initial studies by [63, 29] employed techniques like PHOG and SIFT keypoints for character retrieval. Iwata et al. [137], inspired by [138], adapted these methods specifically for manga characters, considering characters rather than entire page ROIs. Authors further explored applications of these techniques in [139]. However, recently, Li et al. [5] noted the limitations of handcrafted features, particularly their lack of robustness to variations in pose, expression, and character shape.
In [5], authors tackled manga character retrieval and verification. While retrieval is a known task 8.1, verification refers to the task of classifying whether two images (a pair) belong to the same character or not. These two tasks are challenging since the manga character images have a long-tailed distribution and large quality variations. While addressing challenges posed by long-tailed distribution and quality variations in manga characters, they proposed a novel dual loss approach, combining dual ring loss and dual adaptive re-weighting loss, to overcome biases introduced by imbalanced training data.
In animated cartoons, where characters undergo shape and color changes, [140] introduced an approach based on Self-Supervised Learning (SSL) and Multi-object Tracking (MOT) to identify character clusters and link character proposals, facilitating character re-identification across different scenes.
Zhang et al. [141] identified challenges unique to manga, such as similarities in faces but differences in bodies. They proposed the Face-body and Spatial-temporal Associated Clustering method (FSAC), which uses face-body graphs and spatial-temporal relationship correction to refine clustering and overcome artistic exaggerations and deformations, through the design of a temporal-spatial-related triplet loss to fine-tune the clustering.
Recognizing the challenges of occlusion in manga, Zhang et al. [142] developed Occlusion-Aware Manga Character Re-identification (OAM-ReID) with self-paced contrastive learning. By synthesizing data with occluded speech balloons and incomplete bodies, their framework becomes more effective in learning representative features of characters. They introduce a combination of contrastive, occlusion, and re-identification losses, training a vanilla ViT and TransReID as baselines on 10 Manga109 books.
Soykan et al. [143] leveraged SimCLR [144] with various data augmentation techniques for character re-identification, calling it “Identity-aware SimCLR”, focusing on both local (face) and global (body) representations. Similarly, Sachdeva et al. [41] proposed a multi-task Transformer model (Relationformer [145]) for character clustering and linking detected character bounding boxes at the page level.
Finally, [146] introduced an improved YOLOv5 and XGBoosting approach for object detection, achieving state-of-the-art results across multiple datasets (StyleObject7K, ClipArt1K, Watercolor2K, and Comic2K). Their method highlights the effectiveness of well-annotated training data and extensive data augmentation in detection tasks.
7.1.3 Sentence-based Grounding
@@
In image and video Multimedia processing various works tackle the task with custom CNNs and LSTMs [147] approaches, in an unsupervised manner [148] or with custom multimodal fusion approaches [149, 40].
However, the only work that used a similar approach (GroundingDINO) in a zero-shot setting is [41] which used it to automatically annotate the training set.
7.2 Analysis
Analysis tasks involve a deeper examination of the relationship between text and visual elements in comics, including character-text association, panel sorting, dialog transcription, and translation.
7.2.1 Text-character association
@@
Rigaud et al. [150] tackled this task by using geometric graph analysis and anchor point selection. They developed a distance-based algorithm, both global (across the page) and single-frame (within a panel), relying on Euclidean distances and balloon tail information.
Advancing from Rigaud et al.’s work, Nguyen et al. [151] employed a multitask learning approach with a custom Mask R-CNN (adding a new “PairPool” head for the association prediction) which they called ComicMTL. Their method simultaneously addressed panel and character detection, balloon segmentation, text recognition, and text-speaker association, demonstrating its effectiveness on the DCM772 and eBDtheque datasets.
In [152], authors introduced a novel perspective by considering the continuity of comics. Their work focused on estimating the reading order of frames and texts, vital for associating text with off-panel speakers. However, their algorithmic approach raised concerns regarding scalability and generalizability.
Li et al. [73] extended the Manga109 dataset with the Manga109 dialog dataset, enabling more precise speaker detection. They classified text-character links into “easy” (character in the panel) and “hard” (character not in the panel) configurations, using Faster R-CNN for detection and subsequent association classification. The challenges in this approach, when balloons are spoken alternatively, highlight the potential role of NLP models in handling complex dialogue scenarios.
In a recent study, Sachdeva et al. [41] proposed a multi-task Transformer-based model (Relationformer [145]) with specialized MLP heads for detection and linking tasks. While their model offers an “end-to-end” solution, it operates in two iterations: first for detection and then for linking characters and text. However, their model’s comparison with state-of-the-art methods, particularly in detection, is limited, highlighting the need for more comprehensive benchmarking. In fact, the model does not compare with previous state-of-the-art models (YOLO-based) for detection, but only with publicly available GroundingDino [40], in a zero-shot setting.
The character-text association is a critical but challenging task in comics analysis, requiring a nuanced understanding of narrative structure and visual-textual interplay. Future research may benefit from integrating advanced NLP techniques to better interpret complex dialogue scenarios. Additionally, there is a growing need for unified benchmarks and replicable methodologies, as indicated by the diverse approaches in the LoCU framework, to facilitate comparative analysis and further advancements in this field.
7.2.2 Panel Sorting
@@
Ueno et al. [153] conducted a study revealing that while people can relatively easily identify the first and last panels in a sequence, accuracy diminishes with the presentation of all panels from four-panel strips. This suggests the complexity involved in discerning narrative continuity in comics.
The same study attempted to automate panel sorting using an AlexNet-like architecture. However, the CNN’s performance was poor, indicating the difficulty of the task and the need for more advanced or specialized algorithms in panel sorting.
7.2.3 Dialog transcription
@@
The process typically involves several steps: (i) detecting panels, text boxes, and character boxes, (ii) associating text with speakers, and (iii) sorting text boxes according to their reading order to generate a coherent dialogue transcript.
This task has been addressed by [73] for Japanese Mangas, proposing the needed annotations for Manga109, and more recently by [41] for English Mangas called PopManga. However, a few limitations arise from these approaches: sorting the text boxes is always done as a postprocessing operation, depending only on algorithmic approaches. Cases in which balloons alternate among close panels are inherently mistaken. Recently, Sachdeva et al. [43] proposed Magiv2 which tackled the problem of dialog transcription adding also names (from a name bank) allowing it to operate on the whole collection of pages.
In CoMix [42], a first benchmark for this task was proposed, authors enhanced the previous annotations and proposed a new metric for character naming (based on ANLS) and dialog transcription (based on Hungarian matching and edit distance).
7.3 Segmentation
Segmentation in comic analysis, particularly instance segmentation, presents a nuanced approach to identifying and isolating various elements within a comic panel.
7.3.1 Instance segmentation
@@
By focusing on segmentation instead of detection, we can achieve more accurate panel separation and representation. This precision is especially beneficial when reconfiguring or reformatting comics for different layouts. It ensures that each panel is cleanly isolated, without the remnants of adjacent panels or the original layout, which is a common issue with less precise bounding boxes.
One of the key challenges in instance segmentation within comics is dealing with the diversity of artistic styles and the complexity of layouts. Comics often feature intricate designs, overlapping elements, and varying levels of detail, which can complicate the segmentation process. As described in [47], there are dozens of configurations in which panels appear to respect each other (pure grid, vertical and horizontal staggering, bleed, inset, overlap, etc.). Addressing the task of separating panels and understanding their relative positions within Cohn nomenclature is hard with bounding boxes when delimitations are not precise.
Future research in comic instance segmentation could explore advanced deep-learning models that are adept at handling these complexities. The development of specialized algorithms that can adapt to the unique characteristics of different comic styles would significantly advance the field. There’s also potential for integrating instance segmentation with other tasks like character-text association and panel sorting to create a more comprehensive comic analysis system.
Across the presented datasets, no one proposes semantic segmentations.
7.4 Future Tasks
The Layer of Comics Understanding is a general framework with which we aim to collect various Vision-Language tasks. Therefore, it includes also tasks that are currently not explored in comics, such as Translation.
7.4.1 Translation
Definition: The task of translation in comics refers to translating the textual elements that appear in a panel or a page, according to the scene. It goes beyond simple text translation, intertwining linguistic, graphical, and spatial challenges. It necessitates an integrated approach that respects the visual and narrative structure of the comic.
While translation in comics may seem like a straightforward NLP problem, it involves unique complications. Text in comics appears in various forms such as balloons, onomatopoeias, and scene text, each requiring a different approach. In Balloon Text, translated text must fit within the space of the original speech balloon, preserving the visual integrity of the panel. Complications arise when the text to be translated contains elements like sarcasm, irony, jokes, or idiomatic expressions. In Onomatopoeias, both linguistic translation and graphic replacement are often required to maintain the style consistent with the target language. Lastly, in Scene Text, translating involves both linguistic translation and potentially replacing the text within the scene, which can be a complex graphic design task.
8 Layer 3: Retrieval and Modification
Layer 3 in the Layers of Comics Understanding focuses into advanced image and text understanding through tasks such as Retrieval (uni-modal and multi-modal) and Modification.
8.1 Retrieval
8.1.1 Unimodal Retrieval
@@
Pioneering works in this area have explored various aspects of comic retrieval. Wu et al. [154] developed a comic retrieval system that leveraged OCR and external text.
One significant article [138] focused on retrieving full pages or books containing similar elements. This approach is particularly relevant in copyright violation contexts. The authors implemented feature vectors for each image, storing them in a database, and used these vectors to find the most similar documents. Their feature representation method was based on the Histogram of Oriented Gradients (HOG), which outperformed the traditional SIFT method. Notably, they applied HOG not to the entire images but selectively to Regions Of Interest (ROIs) identified in the comics, such as faces or specific markings.
Building on this approach, authors in [137] utilized a similar methodology for manga character retrieval. They focused on identifying and extracting distinctive features from characters, enhancing the ability to retrieve specific characters from large databases.
8.1.2 Cross-modal retrieval
@@
Sketch-based Retrieval. Matsui et al. [32] introduced a framework for retrieving manga images from sketches using the Fine Multi-scale Edge Orientation Histogram (FMEOH). This method indexed different-sized squares on a page and enabled efficient retrieval from sketch queries. Further advancing this area, authors [155] proposed a two-step system that first processes manga and then retrieves images based on sketches. This framework was one of the first to employ the Manga109 dataset and included margin labeling, objectness-based edge orientation histograms, and approximate nearest neighbor search. Narita et al. [156] explored CNN-based feature extraction for sketches and manga images (without and with screentone, respectively), demonstrating improvements over algorithmic approaches.
@@
Multimodal Learning and Character Retrieval. Nguyen et al. [157] proposed a multitask multimodal approach for character image and text learning, aiming at tasks such as character retrieval and emotion recognition. Their method focused on leveraging both visual and verbal information in manga content, in a CLIP-like style aiming at performing various character tasks (e.g. detection, retrieval, emotion classification), trained uniquely with visual information and verbal information in manga image content. Later, in [158] authors introduced ComicLib, a dataset for comic sketch research, providing benchmarks across various tasks like colorization, generation, and retrieval. They conducted extensive comparison experiments with other datasets (e.g. QuickDraw [159]) to provide a benchmark for ComicLib on common tasks like colorization (Style2Paints [160]), generation (DCGAN [161]), retrieval (image-to-image with ResNet), detection (with Faster R-CNN [162], YOLOv3 [163], and SSD [164]) and recognition. However, the dataset’s availability is unclear.
@@
Text-to-Image Retrieval. Shen et al. [165] presented a system for text-to-image content retrieval within Manga frames. This multi-staged system integrates object detection, text recognition, and a vision-text encoder to facilitate efficient search and retrieval of dialogues and scenes. Their method is composed of an object detection model for identifying text and frame bounding boxes (DETR-based [166]), a Vision Encoder-Decoder model for text recognition (DiT [167]), a text encoder for embedding text (multilingual DistilBERT [168]), and a vision-text encoder with unified embedding space (CLIP [14]). They experimented with Japanese and GPT-4101010Using the website https://platform.openai.com/translate/rephrased English sentences, on the annotated panels from “Dollgun Book”. Despite being a composition of existing systems on panel level, on a small distribution of data, the work is a first direction to more complex Comic retrieval.
8.2 Modification

### v2

@@ 7.1.2. Character Re-Identification
Initial studies by Khan et al. (Khan et al., 2012) and Amirshahi et al. (Amirshahi et al., 2012) employed techniques like PHOG and SIFT keypoints for character retrieval. Iwata et al. (Iwata et al., 2014), inspired by Sun et al. (Sun and Kise, 2011), adapted these methods for manga characters, focusing on character-level identification rather than entire page regions, and further explored their applications in (Iwata et al., 2015). However, Li et al. (Li et al., 2021) highlighted the limitations of these handcrafted features, particularly their sensitivity to pose, expression, and shape variations. Li et al. (Li et al., 2021) addressed manga character retrieval and verification, the latter being the task of classifying whether two images belong to the same character. They introduced a dual loss approach—combining dual ring loss and dual adaptive re-weighting loss—to address long-tailed distributions and quality variations in manga data, improving robustness. In animated cartoons, where characters frequently change shape and color, Nir et al. (Nir et al., 2022) proposed a Self-Supervised Learning (SSL) and Multi-object Tracking (MOT) approach to cluster and link character proposals, facilitating character re-identification across scenes. Similarly, Zhang et al. (Zhang et al., 2022) tackled the challenges of character re-identification in manga, where faces are similar but body features differ. They introduced FSAC (Face-body and Spatial-temporal Associated Clustering), which refines clustering through face-body graphs and spatial-temporal relationships, using a triplet loss to handle artistic exaggerations. To address occlusion in manga, Zhang et al. (Zhang and Chu, 2023) developed Occlusion-Aware Manga Character Re-identification (OAM-ReID), using self-paced contrastive learning. Their model synthesizes occluded data, such as speech balloons and incomplete bodies, and combines contrastive, occlusion, and re-identification losses, training a ViT and TransReID on 10 Manga109 books. Soykan et al. (Soykan et al., 2023) introduced ”Identity-aware SimCLR” for character re-identification, leveraging SimCLR (Chen et al., 2020) with data augmentations to capture both local (face) and global (body) features. Similarly, Sachdeva et al. (Sachdeva and Zisserman, 2024) proposed Relationformer (Shit et al., 2022), a multi-task Transformer model for character clustering and linking bounding boxes at the page level. Finally, Ahmad et al. (Ahmad and Schich, 2023) introduced an enhanced YOLOv5 combined with XGBoost for object detection, achieving state-of-the-art performance across multiple datasets, emphasizing the importance of well-annotated data and augmentation techniques in improving detection tasks.
7.1.3. Sentence-based Grounding
@@
In image and video Multimedia processing various works tackle the task with custom CNNs and LSTMs (Li et al., 2019) approaches, in an unsupervised manner (Parcalabescu and Frank, 2020) or with custom multimodal fusion approaches (Shen et al., 2023b; Liu et al., 2023c). However, the only work that used a similar approach (GroundingDINO) in a zero-shot setting is (Sachdeva and Zisserman, 2024) which used it to automatically annotate the training set.
7.2. Analysis
Figure 9. Illustration of Analysis tasks.
Analysis tasks involve a deeper examination of the relationship between text and visual elements in comics, including character-text association, panel sorting, dialog transcription, and translation (Fig. 9).
7.2.1. Text-character association
@@
Rigaud et al. (Rigaud et al., 2015) addressed the text-character association task using geometric graph analysis and anchor point selection, developing a distance-based algorithm that utilized both global (page-level) and single-frame (panel-level) information, relying on Euclidean distances and balloon tail cues. Building on this, Nguyen et al. (Nguyen et al., 2019a) introduced ComicMTL, a multitask learning approach based on a modified Mask R-CNN with an added ”PairPool” head for association prediction. Their method tackled multiple tasks—panel and character detection, balloon segmentation, text recognition, and text-speaker association—demonstrating effectiveness on DCM772 and eBDtheque datasets. Omori et al. (Omori et al., 2022) took a novel approach by focusing on the continuity of comics, particularly reading order, to associate text with off-panel speakers. However, their approach raised scalability and generalizability concerns. Li et al. (Li et al., 2023a) expanded the Manga109 dataset by introducing the Manga109 Dialog dataset, refining speaker detection. They categorized text-character links into “easy” (character in the panel) and “hard” (character off-panel), employing Faster R-CNN for detection and association classification. Despite progress, handling alternating speaker dialogue remains challenging, pointing to the potential of NLP models for complex scenarios. Sachdeva et al. (Sachdeva and Zisserman, 2024) recently proposed a multi-task Transformer model, Relationformer (Shit et al., 2022), with specialized MLP heads for detection and linking. Their two-step, end-to-end solution first detects and then links characters and text, though their model’s benchmarking, especially against YOLO-based models, remains limited, focusing only on GroundingDino (Liu et al., 2023c) in a zero-shot setting. Lastly, in (Soykan et al., 2024b) built on top of Comic-MTL (Nguyen et al., 2021a) a Mask R-CNN based model for detecting and linking textboxes and characters, but did not fully compared with previous models.
Character-text association is a crucial yet challenging task in comics analysis, requiring an understanding of narrative and visual-textual interplay. Future research could benefit from advanced NLP techniques to manage complex dialogue scenarios, and more unified benchmarks and replicable methodologies are needed, as highlighted by the LoCU framework, to support comparative analysis and innovation in this field.
7.2.2. Panel Sorting
@@
Ueno et al. (Ueno and Isahara, 2017) conducted a study revealing that while people can relatively easily identify the first and last panels in a sequence, accuracy diminishes with the presentation of all panels from four-panel strips. This suggests the complexity involved in discerning narrative continuity in comics. The same study attempted to automate panel sorting using an AlexNet-like architecture. However, the CNN’s performance was poor, indicating the difficulty of the task and the need for more advanced or specialized algorithms in panel sorting.
7.2.3. Dialog transcription
@@
The process typically involves several steps: (i) detecting panels, text boxes, and character boxes, (ii) associating text with speakers, and (iii) sorting text boxes according to their reading order to generate a coherent dialogue transcript. This task has been addressed by (Li et al., 2023a) for Japanese Mangas, proposing the needed annotations for Manga109, and more recently by (Sachdeva and Zisserman, 2024) for English Mangas called PopManga. However, a few limitations arise from these approaches: sorting the text boxes is always done as a postprocessing operation, depending only on algorithmic approaches. Cases in which balloons alternate among close panels are inherently mistaken. Recently, Sachdeva et al. (Sachdeva et al., 2024) proposed Magiv2 which tackled the problem of dialog transcription adding also names (from a name bank) allowing it to operate on the whole collection of pages. In CoMix (Vivoli et al., 2024a), a first benchmark for this task was proposed, authors enhanced the previous annotations and proposed a new metric for character naming (based on ANLS) and dialog transcription (based on Hungarian matching and edit distance).
7.2.4. Translation
Definition: The task of translation in comics involves not only linguistic challenges but also graphical and spatial ones, requiring an integrated approach that respects the visual and narrative structure of the comic. Text appears in various forms such as balloons, onomatopoeias, and scene text, each necessitating different strategies. For instance, Balloon Text must fit within the original speech balloon, preserving visual integrity, while Onomatopoeias may require both linguistic translation and graphic replacement to maintain stylistic consistency.
Recently, several works have addressed these complexities by integrating multimodal information and context-aware techniques. (Hinami et al., 2020) were among the first to incorporate context obtained from manga images into a multimodal translation framework, developing a comprehensive system for fully automated manga translation. To improve text image translation performance, (Ma et al., 2023) proposed a Multi-Teacher Knowledge Distillation method, effectively distilling knowledge from a pipeline model into an end-to-end Text Image Machine Translation (TIMT) model. The advent of Large Language Models (LLMs) has further advanced the field; (Yang et al., [n. d.]) explored leveraging LLMs for manga translation tasks, demonstrating the potential of these models in handling complex scenarios involving both text and visual elements. Similarly, (Lippmann et al., 2024) proposed a methodology that leverages the vision component of multimodal LLMs to improve translation quality, investigating the impact of translation unit size and context length, and introducing a token-efficient approach along with a new evaluation dataset. Inspired by human translation workflows, (Singh et al., 2024) presented an approach to completely automate the process of translating graphic novels, addressing the challenges of automating different aspects of the translation workflow. Focusing on character identification and speaker prediction, (Li et al., 2024a) proposed an iterative multimodal framework that employs multimodal information for both tasks, allowing machines to identify characters and predict speaker names based solely on unannotated comic images. Furthermore, (Liu et al., 2024) introduced the task of character-centric story generation, presenting the first model capable of generating visual stories with consistently grounded and coreferent character mentions, emphasizing the importance of character consistency and coreference in narrative understanding and translation. Despite these tasks being tackled as image-to-text translation, they have origins in image-to-image translation due to possible changes in visual components influenced by cultural elements, as discussed by (Khanuja et al., 2024). This indicates that future research may focus more on end-to-end image-to-image translation and transcreation, where both textual and visual elements are adapted to enhance cultural relevance.
7.3. Segmentation
Figure 10. Illustration of Segmentation tasks.
Segmentation in comic analysis, particularly instance segmentation, presents a nuanced approach to identifying and isolating various elements within a comic panel (Fig. 10).
7.3.1. Instance segmentation
@@
Instance segmentation, as opposed to detection, offers more accurate panel separation and representation in comics, which is crucial when reconfiguring or reformatting them for different layouts. This precision ensures that each panel is cleanly isolated, avoiding remnants of adjacent panels or the original layout—a common issue with less precise bounding boxes. One of the key challenges in instance segmentation for comics lies in the diversity of artistic styles and the complexity of layouts. Comics often feature intricate designs, overlapping elements, and various configurations such as grids, staggered layouts, bleeds, and insets, as noted by Cohn (Cohn, 2023). These complex panel arrangements make it difficult to segment panels accurately using bounding boxes, especially when panel boundaries are not clearly defined. In a recent work (Kouletou et al., 2024), authors proposed a pipeline of detection models (YOLO) and segmentations (GroundingDINO) for segmenting objects in comic panels. Despite the idea being simple and elegant -using in a cascade approach a YOLO detector and prompting SAM with the detected bounding boxes- the approach has been only evaluated for the detection tasks as no Segmentation ground truth exists for the used datasets. Moreover, as noticed in (Vivoli et al., 2024c), results are not comparable due to different evaluation datasets/splits/criteria.
Although current datasets do not include semantic segmentations, the 2024 “AI for Visual Art” workshop (Bhattacharjee and Aydemir, 2024) introduced a challenge on comic semantic segmentation, signaling the growing importance of this task in comic analysis.
8. Layer 3: Retrieval and Modification
Layer 3 in the Layers of Comics Understanding focuses on advanced image and text understanding through tasks such as Retrieval (uni-modal and multi-modal) and Modification.
8.1. Retrieval
Figure 11. Illustration of Retrieval tasks.
The task of retrieval, in the domain of the comic, addresses two main retrieval settings in: unimodal and cross-modal retrieval, which do not differ, in practice, from the vision-language retrieval tasks (Fig. 11). These tasks aim to retrieve an instance of the same modality of the query (unimodal) or a different modality from the query (cross-modal), using either text or images as the query. An additional task emerged recently in the realm of Vision-Language, called “Composed image retrieval”, which combines in the query both image and text.
8.1.1. Unimodal Retrieval
@@
Pioneering works in this area have explored various aspects of comic retrieval. Wu et al. (Wu, 2010) developed a comic retrieval system that leveraged OCR and external text. One significant article (Sun and Kise, 2011) focused on retrieving full pages or books containing similar elements. This approach is particularly relevant in copyright violation contexts. The authors implemented feature vectors for each image, storing them in a database, and used these vectors to find the most similar documents. Their feature representation method was based on the Histogram of Oriented Gradients (HOG), which outperformed the traditional SIFT method. Notably, they applied HOG not to the entire images but selectively to Regions Of Interest (ROIs) identified in the comics, such as faces or specific markings. Building on this approach, authors in (Iwata et al., 2014) utilized a similar methodology for manga character retrieval. They focused on identifying and extracting distinctive features from characters, enhancing the ability to retrieve specific characters from large databases.
8.1.2. Cross-modal retrieval
@@
Sketch-based Retrieval. Matsui et al. (Matsui et al., 2014) introduced a framework for retrieving manga images from sketches using the Fine Multi-scale Edge Orientation Histogram (FMEOH). This method indexed different-sized squares on a page and enabled efficient retrieval from sketch queries. Further advancing this area, authors (Matsui et al., 2015) proposed a two-step system that first processes manga and then retrieves images based on sketches. This framework was one of the first to employ the Manga109 dataset and included margin labeling, objectness-based edge orientation histograms, and approximate nearest neighbor search. Narita et al. (Narita et al., 2017) explored CNN-based feature extraction for sketches and manga images (without and with screentone, respectively), demonstrating improvements over algorithmic approaches.
@@
Multimodal Learning and Character Retrieval. Nguyen et al. (Nguyen et al., 2021a) proposed a multitask multimodal approach for character image and text learning, aiming at tasks such as character retrieval and emotion recognition. Their method focused on leveraging both visual and verbal information in manga content, in a CLIP-like style aiming at performing various character tasks (e.g. detection, retrieval, emotion classification), trained uniquely with visual information and verbal information in manga image content. Later, in (Wei et al., 2022) authors introduced ComicLib, a dataset for comic sketch research, providing benchmarks across various tasks like colorization, generation, and retrieval. They conducted extensive comparison experiments with other datasets (e.g. QuickDraw (Ha and Eck, 2017)) to provide a benchmark for ComicLib on common tasks like colorization (Style2Paints (Zhang et al., 2017)), generation (DCGAN (Radford et al., 2015)), retrieval (image-to-image with ResNet), detection (with Faster R-CNN (Ren et al., 2015), YOLOv3 (Redmon et al., 2015), and SSD (Liu et al., 2015a)) and recognition. However, the dataset’s availability is unclear.
@@
Text-to-Image Retrieval. Shen et al. (Shen et al., 2023a) presented a system for text-to-image content retrieval within Manga frames. This multi-staged system integrates object detection, text recognition, and a vision-text encoder to facilitate efficient search and retrieval of dialogues and scenes. Their method is composed of an object detection model for identifying text and frame bounding boxes (DETR-based (Carion et al., 2020)), a Vision Encoder-Decoder model for text recognition (DiT (Touvron et al., 2020)), a text encoder for embedding text (multilingual DistilBERT (Sanh et al., 2019)), and a vision-text encoder with unified embedding space (CLIP (Radford et al., 2021)). They experimented with Japanese and GPT-4888Using the website https://platform.openai.com/translate/rephrased English sentences, on the annotated panels from “Dollgun Book”. Despite being a composition of existing systems on panel level, on a small distribution of data, the work is a first direction to more complex Comic retrieval.
8.2. Modification
Figure 12. Illustration of Modification tasks.